# Self-Controlled Diffusion for Denoising in Scientific Imaging

### v1

@@ Abstract
This paper presents a novel approach for denoising Electron Backscatter Diffraction (EBSD) patterns using diffusion models. We propose a two-stage training process with a UNet-based architecture, incorporating an auxiliary regression head to predict the quality of the experimental pattern and assess the progress of the denoising process. The model uses an adaptive denoising strategy, which integrates quality prediction and feedback-driven iterative denoising process control. This adaptive feedback loop allows the model to adjust its schedule, providing fine control over the denoising process. Furthermore, our model can identify samples where no meaningful signal is present, thereby reducing the risk of hallucinations. We demonstrate the successful application of diffusion models to EBSD pattern denoising using a custom-collected dataset of EBSD patterns, their corresponding master patterns, and quality values.
@@ 1 Introduction
EBSD, or Electron Backscatter Diffraction, is a technique used in electron microscopy to study the crystallographic properties of materials at the nanoscale. In this method, electrons from an electron beam are diffracted by crystallographic planes, creating a diffraction pattern that can be analyzed to determine the crystal orientation. Due to its high resolution, EBSD can provide detailed information about grains, including those that are nanometer-sized.
@@
Figure 1: a) Pattern formation as a tangential projection of diffraction sphere on a detector plane; b) Inverse Polar Figure (IPF) map visualising the orientation of a grain by the color of the crystal face pointing perpendicular to the image plane. Each point of IPF map is the result of EBSD pattern indexing.
@@
EBSD patterns appear as grayscale images featuring diffraction lines, known as Kikuchi lines. Each EBSD pattern represents a portion of the entire diffraction sphere. Similar to maritime navigation, where specific sections of the sky help determine one’s location, the Kikuchi pattern is utilized to identify the crystallographic orientation of individual crystallites (Figure 1). Process of determination of crystallographic orientation by analysis of Kikuchi patterns is called indexing (assigning crystallographic indexes to every Kikuchi line appear at the pattern).
@@
The quality of EBSD results depends on several parameters, including those related to the electron microscope (such as probe current, accelerating voltage, working distance, and magnification), the type of detector (e.g., CCD or single electron detectors), and, most importantly, the sample itself (surface roughness, composition, crystallization degree, chemical and thermal stability and so on) [18, 7]. Stable and well-crystallized inorganic materials, like steel and its alloys or alumina, produce clear and distinguishable Kikuchi patterns [4, 3]. In contrast, some beam-sensitive materials, such as perovskite which become now the most promising materials for the green energy transition as active layer material in solar cells, or certain metals in thin film form, generate weak and noisy patterns [16, 7].
The analysis of weak patterns is a complex task. Currently, there are three main approaches to patterns indexing: Hough indexing (HI) [8, 15], Dictionary Indexing (DI) [20] and relatively recently developed Spherical indexing (SI) [17].
Hough Indexing converts an EBSD pattern into Hough space to summarise the intensity of the entire Kikuchi line into a single point. This representation allows for the detection of that line in the original pattern. By analysing the overall position of all lines, the crystal orientation can be determined. However, a major drawback of this method is its requirement for high-quality patterns, which limits its applicability to beam-sensitive materials.
The Dictionary Indexing (DI), on the other hand, offers much greater precision in determining orientation, even for somewhat noisy patterns. This method involves using an algorithm that compares complete experimental EBSD patterns against simulated patterns (Figure 2) [20]. Despite its advantages, the DI is quite slow because it necessitates the generation of a large number of parts of the simulated Master Pattern and massive amount of calculations during analysis.
Currently, the most advanced indexing method is the Spherical Indexing (SI). Similar to the DI, it requires the generation of a complete diffraction picture known as the Master Pattern. However, instead of calculating a dot product for each experimental pattern and piece of the Master Pattern, the Spherical Indexing utilizes spherical functions, resulting in higher precision and much faster analysis. However, it still necessitates considerable processing power.
Figure 2: An example of Master Pattern (MP) and its components identified during Spherical Indexing. The same components of MP served as ground truth in model training.
However, like all indexing techniques, SI’s effectiveness relies heavily on the pattern preprocessing procedure. Adjustments to the image processing steps can lead to significant changes in results, and it is not always clear which procedures should be applied and in what order. Presently, achieving optimal results often depends on experience and, at times, luck, as the signal quality can vary from sample to sample and from measurement to measurement. Additionally, the current capabilities of SI allow for the indexing of highly noisy patterns, where not a single trace of a Kikuchi line is detectable by the naked eye. Therefore, it can be difficult to determine whether the indexing is truly accurate.
In this context, EBSD pattern restoration might be a valuable part of the EBSD post-analysis. The ability to restore noisy patterns opens up possibilities for improved analysis in three key ways:
  • enhance pattern processing to enable indexing of patterns that were previously unindexable;
@@
  • make indexing more reliable, consistent, and less dependent on the preprocessing of patterns.
The last two points can be partially addressed by utilizing advanced detectors, such as single-particle-counting detectors (e.g., Medi/Timepix series), which provide noiseless digital integration (counting) and high detection efficiency [11]. A particle-counting detector functions such that each pixel serves as a detector. Each photon (or electron) that reaches a pixel is compared to a predefined threshold, and a digital counter is incremented accordingly. These detectors are particularly effective when signal intensity is low and collecting substantial signal for adequate statistical noise suppression would damage the sample. Despite their advantages, they are often insufficient for beam-sensitive materials, and the resulting images still require further processing before they can be indexed. Though, combination of them, such as single photon detector and denoising system restoring noisy patterns to a more clear EBSD patterns, can become a very powerful solution helping to reveal new insights about materials and making EBSD method more common and easy in use making it more accessible for more researchers.
Computer vision approaches offer a promising way to denoise EBSD patterns and extract more valuable information from them. Currently, several methods are employed for this purpose: Autoencoders, Generative Adversarial Networks (GANs), and diffusion-based learning techniques. While researchers have applied the first two approaches to denoise EBSD patterns, their effectiveness is constrained by pattern quality, which must remain sufficiently high to ensure stable results and prevent hallucinations during restoration. Diffusion learning offers an advantage through its step-by-step procedural approach. However, as we will demonstrate in the Related Works section, even these sophisticated techniques can produce hallucinations, particularly when processing content-simple images like EBSD patterns.
To address these challenges and improve the reliability of EBSD pattern denoising, this paper aims to develop a control mechanism that limits the restoration process while preserving vital information in grayscale, line-based EBSD images. Our contributions include:
@@ 2 Related Works
Improving the quality of collected EBSD patterns (EBSP) using Deep Learning (DL) approaches is an active topic in recent years. A review of existing literature reveals two main strategies of doing this:
  • Directly enhancing EBSP to improve the quality of standard indexing procedures. This involves collecting EBSP with sequentially improved visibility of Kikuchi lines or an enhanced Signal-to-Noise Ratio (SNR). These DL optimized patterns can then be analyzed using SI/DI/HI for orientation determination;
  • Indexing by SI, DI, or HI methods EBSP that have been processed using traditional imaging techniques. This step is followed by the enhancement of the resulting Inverse Pole Figure (IPF) maps through Deep Learning.
The first approach appears to be the more convenient choice for enhancing the quality of EBSD results. Since the pattern dataset serves as the foundation for the entire process, starting with this step could lead to significant overall improvements in the analysis.
@@
The initial studies on denoising EBSD patterns utilized UNet-based autoencoders [2] where the authors achieved a commendable quality of restored patterns. At that time, Spherical Indexing was not widely adopted, and the initial patterns they worked with were often poorly indexed using Hough Indexing. Nowadays, SI can effortlessly index patterns of such quality. Consequently, a significant portion of this work is focused on enhancing the precision of strain analysis, which requires much higher-quality patterns.
It’s important to highlight that there are very few papers dedicated to pattern restoration when the patterns are predominantly noise and nearly devoid of detectable lines to the human eye. One notable exception is the work by [12], where the patterns are nearly lost, although an experienced user can still identify some Kikuchi lines. The lack of research addressing such low-quality patterns arises from the lack of methods to verify whether predictions are accurate and the challenges associated with collecting ground truth data.
The second approach is less common and requires more effort in data collection, as an Inverse Pole Figure (IPF) map consists of thousands of EBSD patterns. To obtain reliable results, a large number of experiments must be conducted to create a sufficiently sized dataset based on IPF maps. This process often involves significant data augmentation, which can lead to the sacrifice of some details within the data. In the work by [14], the IPF maps were restored using the Diffusion approach. To stabilize the diffusion model, they utilized a Reduced-Order Generalized Spherical Harmonic (ROGSH), which incorporates additional information to enhance restoration performance.
Diffusion models have become one of the most powerful approaches for image restoration, particularly for denoising, due to their ability to restore images gradually and more carefully compared to alternatives such as GANs or VAEs. However, this stability comes at the cost of significant computational resources needed for the restoration process that might be critical in terms of using such networks for ”every-day” EBSD dataset restoration what usually contains thousands or even tens of thousand images.
@@
Nevertheless, the main advancement of diffusion models that enhance image restoration precision is their approach to handling noisy images. Unlike classical models such as GANs or VAEs, which aim to transform a noisy sample directly into a denoised one, diffusion models transform a noisy image into a less noisy version in several steps which is especially valuable in application to scientific images which are formed due to physical properties of samples. This gradual restoration occurs within the space of images, transitioning from a noisy state to a denoised one. This approach was originally proposed not just for diffusion models, but as a more general method for restoring images without the need to train separate models for every restoration problem [10].
@@
The diffusion models are represented by several types of approaches. The Denoising Diffusion Probabilistic Model introduced by [5] uses Markov chains to model the forward and reverse process. Step by step it applies noise (usually Gaussian) during forward process. Then in backward process the model is learning noise distribution parameters such as mean $\mu_{\theta}(x_{t},t)$ and variance $\beta_{\theta}(x_{t},t)$ to be able to restore image from noisy state back to original state.
@@
Despite their popularity in many imaging tasks recently, the diffusion models are still not widely used for scientific image and especially EBSD images denoising. However, it proved its efficiency in denoising regular images like photographs [19]. The crucial difference in the [19] work is in method of sampling they applied. In previous works, every next image in the denoising pipeline is a combination of restored image from the previous state and normalized noise which is left after subtraction of restored image from the noisy one (1). In the [19] work, every next image is an image from the previous step and subtracted normalized difference between original noisy image and a restored image from the previous step (2). Changing the sampling method has helped stabilize the image restoration process.
$x_{t-1}=S_{\theta}(x_{t},t)+\frac{(t-1)}{T}\cdot[x_{T}-S_{\theta}(x_{t},t)]$ (1)
$x_{t-1}=x_{t}-\frac{1}{T}\cdot[x_{T}-S_{\theta}(x_{t},t)]$ (2)
@@ 3 Dataset
There are EBSD datasets available on the Internet (https://zenodo.org/communities/ebsd/ ). However, these datasets were not used in this work as our approach relies on Master Patterns as a ground truth and employs a patterns quality metric, which is not available in the public data. Therefore, we collected new datasets and generated corresponding pieces of Master Patterns using Spherical Indexing.
To collect the experimental dataset for model training, we utilized a Scanning Electron Microscope Verios 460 from ThermoFisher, equipped with an EDAX Clarity detector utilizing a 4-chip Timepix detector from Amsterdam Scientific Instruments (ASI). The data patterns were gathered using EDAX APEX software with beam currents ranging from 50 to 200 pA and accelerating voltages between 5 and 20 kV. Pixel integration times varied from 30 to 250 ms, with different binning options applied (1, 2, 4, and 5 pixels). We analyzed multiple samples (Nickel ($Ni$), Ruthenium ($Ru$), Gold ($Au$), Silver ($Ag$)) to apply a diverse set of measurement conditions and to include various crystallographic variables, including different crystal lattice parameters and crystal systems (Cubic, Orthorhombic, Hexagonal). In total, the number of experimental patterns used for training was over 163000, with further 16500 patterns used for validation and over 6000 used as a hold-out test set.
@@ 3.1 Master patterns
Master patterns for each crystal structure were generated using EDAX OIM 9 software. This software was also utilized to index experimental EBSD patterns using Spherical Indexing. The master patterns identified through this process were exported and used as ground truth for training purposes.
@@ 4.2 Training procedure
Denoising model. On the first stage we training the main UNet-style encoder-decoder denoising neural network to perform image denoising. The pseudocode for this training procedure is presented in Algorithm 1, which is depicted in Figure 3(a). During training, the network receives a noisy observation $x_{t}$ as input and attempts to predict the corresponding clean image $x_{0}$. The denoising process is guided by a quality score $q$ and a noise scale $s_{\text{noise}}$, with additional Gaussian noise injected at each diffusion step to simulate varying levels of corruption.
The noise injection process follows a diffusion model, where the noisy observation $x_{t}$ is computed as a weighted combination of the clean image $x_{0}$, the noisy input $x$, and an additional noise term $\omega_{t}$ drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$, Eq. 1.
@@
Notably, the algorithm allows the generation of images that can be more degraded than the observed image $x$, since for each experimental image $t_{x}\leq T$ and the current step $t$ can exceed $t_{x}$, leading to further simulated corruption of the image. This, together with the extra noise enables the model to recover clean images from increasingly corrupted data.
On each step, the network has access to the interpolated noisy image of the step $x_{t}$ and is guided by the original experimental noisy image $x$, which gives the network more context (both are stacked as a two channel image), and the current step $t$. The network’s parameters $\theta$ are updated through backpropagation to minimize the loss function $\mathcal{L}(\hat{x}_{0},x_{0})$, which measures the difference between the predicted clean image $\hat{x}_{0}$ and the true clean image $x_{0}$.
@@
Quality Values Regression. In the second stage of training (Algorithm 2, depicted in Figure 3(b)), we perform quality regression using the trained denoising network in a frozen state. Specifically, the bottleneck feature map, which captures an internal representation of the input tensor, is passed through an auxiliary head for the two-values regression task. The goal is to predict two key values: the original image quality $\hat{q}$, and the progress of the denoising process $\hat{r}_{\text{denoise}}$.
The first predicted value, $\hat{q}$, represents the estimated quality of the original image at the current denoising step, corresponding to the same ground-truth value $q$ from the dataset for all diffusion steps. Since the experimental noisy image $x$ is part of the input tensor at each step, the information necessary for predicting this objective value is already present in the input.
The second predicted value, $\hat{r}_{\text{denoise}}$, estimates the progress of the denoising process. This value depends on the current diffusion step $t$ and provides a measure of how far along the diffusion process is in recovering the clean image. More specifically, $\hat{r}_{\text{denoise}}$ estimates how much noise has been removed by the current denoising step, or in other words, it can be viewed as a measure of how close the image $x_{t}$ is to the final clean state.
The target value for $\hat{r}_{\text{denoise}}$ is constructed using the following rule:
@@
When the image is not fully corrupted (i.e., $x$ is not pure noise), the value is inversely proportional to the diffusion step, with lower $t$ values indicating a more advanced stage of denoising. For fully noisy images, the progress is set to $r_{\text{denoise}}=0.0$. The sample $x$ can represent pure noise in several cases: the dataset includes intentionally collected pure noise EBSD patterns, where noise collected as instrumental noise without a sample material installed, or synthetic noise samples generated during training by augmenting the data with pure noise drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$. For all such cases, both regression targets are set to 0.
@@
7:     Predict regression values: $(\hat{q},\hat{r}_{\text{denoise}})\leftarrow g_{\theta}(f_{\theta}^{enc}(x_{t},x,t_{\text{rand}}))$, where $t_{\text{rand}}\sim\mathcal{U}(0,T)$ random diffusion step.

### v2

# DIFFRACT: Diffusion-based Restoration via Adaptive Control and Thresholding for Diffraction Imaging

@@ Abstract
This paper presents a novel approach for denoising Electron Backscatter Diffraction (EBSD) patterns using diffusion models. We propose a two-stage training process with a UNet-based architecture, incorporating an auxiliary regression head to predict the quality of the experimental pattern and assess the progress of the denoising process. The model uses an adaptive denoising strategy, which integrates quality prediction and feedback-driven iterative denoising process control. This adaptive feedback loop allows the model to adjust its schedule, providing fine control over the denoising process. Furthermore, our model can identify samples where no meaningful signal is present, thereby reducing the risk of hallucinations. We demonstrate the DIFFRACT - the successful application of diffusion models to EBSD pattern denoising using a custom-collected dataset of EBSD patterns, their corresponding Master Patterns, and quality values.
@@ 1 Introduction
EBSD (Electron Backscatter Diffraction) is an electron microscopy technique for studying crystallographic properties at the nanoscale. In this method, electrons are diffracted by crystallographic planes, producing patterns that reveal crystal orientation. Its high resolution provides detailed information on grains, even at the nanometer scale, offering insights into processes such as degradation mechanisms, electronic properties, and other material characteristics.
@@
Figure 1: a) Pattern formation as a tangential projection of diffraction sphere on a detector plane; b) Inverse Polar Figure (IPF) map visualising the orientation of a grain by the color of the crystal face pointing perpendicular to the image plane. Each point of the IPF map is the result of EBSD pattern indexing.
@@
EBSD patterns are grayscale images with diffraction lines called Kikuchi lines, each representing a portion of the diffraction sphere. Similar to maritime navigation using constellations, Kikuchi patterns identify the crystallographic orientation of individual crystallites (Figure 1). This process, known as indexing, assigns crystallographic indexes to each Kikuchi line in the pattern.
@@
The quality of EBSD results depends on several factors, including microscope settings (probe current, accelerating voltage, working distance, magnification), detector type (e.g., CCD or single-electron), and, most importantly, the sample itself (surface roughness, composition, crystallization degree, chemical and thermal stability) [6, 23]. Well-crystallized materials like steel or alumina produce clear Kikuchi patterns [16, 15], whereas beam-sensitive materials such as hybrid lead halide perovskites (promising for solar cell active layers) or metals in thin film form often yield weak, noisy patterns [14, 23].
Analyzing weak patterns is challenging, with three main approaches: Hough Indexing (HI) [12, 18], Dictionary Indexing (DI) [7], and the recently developed Spherical Indexing (SI) [19]. HI processes patterns in Hough space but depends on high-quality data, limiting its use for beam-sensitive materials. DI improves precision for noisy patterns by comparing them to simulated Master Patterns (Figure S6) but is computationally intensive. SI also uses a Master Pattern but matches via spherical functions rather than dot products, enabling faster analysis and higher precision than DI.
However, SI’s effectiveness depends on preprocessing, which is subjective, variable, and often requires experience and, at times, luck. While SI can index very noisy patterns lacking visible Kikuchi lines, assessing indexing accuracy remains difficult. In this context, EBSD pattern restoration can enhance post-analysis in three key ways:
@@
  • make indexing more reliable, consistent, and less dependent on the preprocessing of patterns;
  • enhance pattern processing to enable indexing of patterns that were previously unindexable.
Computer vision approaches offer a promising route for denoising EBSD patterns and extracting more information. Current methods include Autoencoders, GANs, CNNs, and diffusion-based techniques. While the first two have been applied to EBSD, their effectiveness depends on sufficiently high pattern quality to ensure stable results and avoid hallucinations. Diffusion learning provides an advantage with its step-by-step approach, but as shown in the Related Works section, even advanced methods can hallucinate when processing content-simple images like EBSD patterns.
To address these challenges and improve the reliability of EBSD pattern denoising, this paper aims to develop a control mechanism that limits the restoration process while preserving vital information in EBSD images. Our contributions include:
@@ 2 Related Works
Image denoising with Deep Learning has been actively developed for over a decade, with most methods based on convolutional architectures. NAFNet is notable as the first nonlinear-activation-free network and has been applied even to medical image denoising, reducing hallucination risk [29, 5, 26]. Another example is the Context Guided Network for Semantic Segmentation (CGNet) [24, 10], which captures both local details and global context through a cascaded gaze mechanism.
Although strong baselines for image denoising, these methods have limitations in scientific imaging, especially microscopy, where long-distance pixel relationships are critical. For images such as EBSD patterns, approaches like GANs, transformers, and diffusion networks (discussed later) are often better suited.
Improving EBSD pattern (EBSP) quality with Deep Learning has been an active research topic in recent years. While diffraction image denoising often adapts techniques from real-world images, it differs due to distinct noise characteristics (less Gaussian, more Poisson) and the unique physics of image formation, involving reciprocal space and greater emphasis on fine details.
Since the pattern dataset forms the foundation for the entire indexing process, most studies first restore EBSD patterns with DL methods, then index the restored patterns and generate IPF maps.
@@
Early EBSD denoising studies used UNet-based autoencoders [2], achieving good restoration quality. At the time, SI was not widely adopted, and many initial patterns were poorly indexed using HI. Today, SI can easily index patterns of that quality, so much of this earlier work focused on improving strain analysis, which demands higher-quality patterns.
Few studies address pattern restoration when images are mostly noise and nearly devoid of visible lines. A notable exception is [17], where patterns are almost lost but experienced users can still spot some Kikuchi lines. The scarcity of such research stems from the difficulty of verifying prediction accuracy and the challenges of collecting ground truth data.
@@
The key advantage of diffusion models for image restoration is their stepwise handling of noisy images. Unlike classical models that map a noisy sample directly to a clean one, diffusion models progressively reduce noise over several steps, particularly valuable for scientific images shaped by physical sample properties. This gradual restoration occurs in image space, transitioning from noisy to clean, and was originally proposed as a general method for image restoration without training separate models for each task [4].
@@
The diffusion models are represented by several types of approaches. The Denoising Diffusion Probabilistic Model introduced by [11] uses Markov chains to model the forward and reverse process. Step by step it applies noise during forward process. Then in backward process the model is learning noise distribution parameters such as mean $\mu_{\theta}(x_{t},t)$ and variance $\beta_{\theta}(x_{t},t)$ to restore image from noisy state back to original state.
@@
Despite their general popularity and proved efficiency in denoising real-world imagery data, like photographs [27], the diffusion models are still not widely used for scientific image and especially EBSD images denoising. The crucial difference in the [27] work is in method of sampling, which has helped stabilize the image restoration process. In previous works, every next image in the denoising pipeline is a combination of restored image from the previous state and normalized noise which is left after subtraction of restored image from the noisy one. In the [27] work, every next image is an image from the previous step and subtracted normalized difference between original noisy image and a restored image from the previous step.
@@
Another approach to improving EBSD results, presented in the literature, involves processing EBSP using traditional imaging techniques, followed by enhancing the resulting Inverse Pole Figure (IPF) maps through Deep Learning. For example, in [3], IPF maps were restored using a diffusion-based approach. To stabilize the diffusion model, the authors employed Reduced-Order Generalized Spherical Harmonics (ROGSH), which incorporate additional information to improve restoration performance. In contrast, the present work focuses on restoring EBSD patterns directly, as this remains the most precise and reliable method for enhancing EBSD analysis.
@@ 3 Dataset
There are EBSD datasets available on the Internet (https://zenodo.org/communities/ebsd). However, these datasets were not used in this work as our approach relies on Master Patterns as a ground truth and employs a patterns quality metric, which is not available in the public data. Therefore, we collected new datasets and generated corresponding pieces of Master Patterns using Spherical Indexing.
We utilized a scanning electron microscope ThermoFisher Verios 460, equipped with an EDAX Clarity detector utilizing a 4-chip Timepix detector offering noiseless digital integration and high detection efficiency [13] from Amsterdam Scientific Instruments (ASI). The data patterns were gathered using EDAX APEX software with beam currents ranging from 50 to 200 pA and accelerating voltages 5-20 kV. Pixel integration times varied from 30 to 250 ms, with different binning options applied (1, 2, 4, and 5 pixels). We analyzed multiple samples (Nickel ($Ni$), Ruthenium ($Ru$), Gold ($Au$), Silver ($Ag$)) to apply a diverse set of measurement conditions and to include various crystallographic variables, including different crystal lattice parameters and crystal systems (Cubic, Hexagonal). In total, the number of experimental patterns used for training was over 163000, with further 16500 patterns used for validation and over 6000 used as a hold-out test set.
@@ 3.1 Master Patterns
Master Patterns for each crystal structure were generated using EDAX OIM 9 software. This software was also utilized to process and index experimental and restored EBSD patterns using Spherical Indexing. The Master Patterns were used as ground truth for the models training and evaluation.
@@ 4.2 Training procedure
Denoising model. On the first stage we training the main model to perform image denoising. The pseudocode for this training procedure is presented in Algorithm 1, which is depicted in Figure 2(a). During training, the network receives a noisy observation $x_{t}$ as input and attempts to predict the corresponding clean image $x_{0}$. The denoising process is guided by a quality score $q$ and a noise scale $s_{\text{noise}}$, with additional Gaussian noise injected at each diffusion step to simulate varying levels of corruption. The noise injection process follows a diffusion model, where the noisy observation $x_{t}$ is computed as a weighted combination of the clean image $x_{0}$, the noisy input $x$, and an additional noise term $\omega_{t}$ drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$, Eq. 1.
@@
Notably, the algorithm allows the generation of images that can be more degraded than the observed image $x$, since for each experimental image the current step $t$ can exceed $t_{x}$, leading to simulated further corruption of the image. This, together with the extra noise enables the model to recover clean images from increasingly corrupted data.
On each step, the network has access to the interpolated noisy image of the step $x_{t}$ and is guided by the original experimental noisy image $x$, which gives the network more context (both are stacked as a two channel image), and the current step $t$.
@@
Quality Values Regression. In the second stage of training (Algorithm 2, depicted in Figure 2(b)), we perform quality regression using the trained denoising network in a frozen state. Specifically, the bottleneck feature map, which captures an internal representation of the input tensor, is passed through an auxiliary head for the two-values regression task. The goal is to predict two values: the original image quality $\hat{q}$, and the progress of the denoising process $\hat{r}_{\text{denoise}}$.
The first predicted value, $\hat{q}$, represents the estimated quality of the original image at the current denoising step, corresponding to the same ground-truth value $q$ from the dataset for all diffusion steps. Since the experimental noisy image $x$ is part of the input tensor at each step, the information necessary for predicting this objective value is present in the input. $q=0$ for pure noise images.
The second predicted value, $\hat{r}_{\text{denoise}}$, estimates the progress of the denoising process. This value depends on the current diffusion step $t$ and provides a measure of how far along the diffusion process is in recovering the clean image. More specifically, $\hat{r}_{\text{denoise}}$ estimates how much noise has been removed by the current denoising step. The target value for $\hat{r}_{\text{denoise}}$ is constructed using the following rule:
@@
When the image is not fully corrupted (i.e., $x$ is not pure noise), the value is inversely proportional to the diffusion step, with lower $t$ values indicating a more advanced stage of denoising; $r_{\text{denoise}}=0.0$ for pure noisy images. The sample $x$ can represent pure noise in two cases: the dataset includes intentionally collected pure noise EBSD patterns, where noise collected as instrumental noise without a sample material installed, or synthetic noise samples generated during training by augmenting the data with pure noise drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$.
@@
7:  Predict regression values: $(\hat{q},\hat{r}_{\text{denoise}})\leftarrow g_{\theta}(f_{\theta}^{enc}(x_{t},x,t_{\text{rand}}))$, where $t_{\text{rand}}\sim\mathcal{U}(0,T)$.